# Neural Point-based Volumetric Avatar: Surface-guided Neural Points for Efficient and Photorealistic Volumetric Head Avatar

### v1

@@ Abstract
Rendering photorealistic and dynamically moving human heads is crucial for ensuring a pleasant and immersive experience in AR/VR and video conferencing applications. However, existing methods often struggle to model challenging facial regions (e.g., mouth interior, eyes, hair/beard), resulting in unrealistic and blurry results. In this paper, we propose Neural Point-based Volumetric Avatar (NPVA), a method that adopts the neural point representation as well as the neural volume rendering process and discards the predefined connectivity and hard correspondence imposed by mesh-based approaches. Specifically, the neural points are strategically constrained around the surface of the target expression via a high-resolution UV displacement map, achieving increased modeling capacity and more accurate control. We introduce three technical innovations to improve the rendering and training efficiency: a patch-wise depth-guided (shading point) sampling strategy, a lightweight radiance decoding process, and a Grid-Error-Patch (GEP) ray sampling strategy during training. By design, our NPVA is better equipped to handle topologically changing regions and thin structures while also ensuring accurate expression control when animating avatars. Experiments conducted on three subjects from the Multiface dataset demonstrate the effectiveness of our designs, outperforming previous state-of-the-art methods, especially in handling challenging facial regions.
Figure 1: Neural Point-based Volumetric Avatar explores point-based neural representation combined with volume rendering, achieving high-fidelity facial animations (images and depth maps) while maintaining efficiency comparable to mesh-based methods (Tab. 2). During training, NPVA can adaptively allocate more points to challenging facial regions, forming a thicker “shell” (i.e., a higher variance of projected distances onto the face in the normal direction) and increasing capacity as needed. In the leftmost image, we show our rendering and the ground truth (GT) side-by-side on the right and left parts, respectively. Observe the close resemblance between our rendering quality and the GT.
1 Introduction
Realizing photorealistic rendering of an animatable human head is a pivotal goal in computer graphics and vision, which has broad applications such as AR/VR communications [8, 28, 19], gaming [37], and remote collaboration [38].
@@ 1 Introduction
However, providing a satisfying and immersive experience in these applications remains immensely challenging due to our innate ability to express and perceive emotions through subtle facial cues [5]. Existing data-driven learning methods often generate noticeable artifacts in the mouth area and unrealistic hair textures [47, 4, 11, 7, 19, 23]. This limitation is primarily attributed to their underlying mesh-based representations, since the predefined mesh has a fixed topology and limited discretization resolution.
@@
To illustrate this limitation, consider the Deep Appearance Model (DAM) [19] that decodes a head mesh and the corresponding view-specific texture UV map. Despite achieving high-quality rendering results for the skin regions, DAM produces conspicuous artifacts in the mouth and hair regions due to inaccurate correspondences across frames (e.g., mouth interior) and the mesh’s inability to model thin structures (e.g., hair). To alleviate these issues, Pixel Codec Avatar (PiCA) [23] proposes the use of neural textures, allowing the subsequent neural renderer to address inaccurate shape estimation and topological inconsistencies (e.g., closed/open mouth), attaining moderately improved facial geometry and renditions for the mouth region. Similarly, Mixture of Volumetric Primitives (MVP) [21] attaches volumetric primitives (predicted by a CNN) to mesh vertices, replacing vertex colors with more flexible volumetric primitives. However, these enhanced “texture”-like representations remain embedded in a predefined topology and tend to produce blurry results if inaccurate correspondences occur.
@@
Therefore, we propose Neural Point-based Volumetric Avatar (NPVA) that leverages highly flexible neural points [1] and versatile neural volume rendering, enabling sharper rendering for both topologically changing geometries (e.g., mouth interior) and translucent thin hair/beard structures.
@@
Efficient rendering and training are also indispensable for practical applications. To this end, we propose three technical innovations. (1) We introduce a novel depth-guided [16] sampling method that incorporates local depth context information (i.e. a patch), achieving more realistic rendering while reducing the rendering time by 10$\sim$ 10$\times$ 10 × compared to the vanilla NeRF. (2) We develop a lightweight radiance decoding process that eliminates unnecessary per-point processing used in [41, 2], significantly improving rendering efficiency (7$\sim$ 7$\times$ 7 ×) and offering better generalization in our dynamic modeling task. (3) Lastly, to speed up training, we propose a novel Grid-Error-Patch (GEP) ray sampling strategy comprising three stages: a uniform grid-sampling stage for rapid initialization of a coarse result, an error-based importance sampling that delves into more challenging regions, and a patch-based stage to impose high-level perceptual image losses.
@@
  • We propose a novel volumetric representation based on neural points that are dynamically allocated around the surface (i.e., the target expression) for animatable head avatar creation. This representation is inherently capable of better handling thin geometry (e.g., hair) and topological changes (e.g., mouth open and close).
@@
  • Experiments on the Multiface dataset show that our approach produces higher-quality images for novel expressions and novel views while being 70$\sim$ 70$\times$ 70 × faster than NeRF.
@@
Figure 2: Overview of Neural Point-based Volumetric Avatar (NPVA). The core of our approach is a neural point-based volumetric representation (middle), with points distributed around the surface of the target expression. This surface is defined by the low-resolution position map $\hat{\bm{G}}_{o}$ with intermediate supervision. A high-resolution displacement map $\hat{\bm{G}}_{d}$ allows the points to adaptively move within a certain range, as needed to provide increased capacity in more challenging regions (e.g., mouth, hair/beard). The attached point features are obtained from the feature map $\hat{\bm{F}}$. $\hat{\bm{G}}_{o}$, $\hat{\bm{G}}_{d}$, and $\hat{\bm{F}}$ are decoded from the latent code $\bm{z}$ (left), which is trained in a variational auto-encoding style (encoder omitted). In addition, we propose three technical innovations with the aim of achieving rendering efficiency on par with mesh-based methods (right).
2 Related Works
@@
Continuous Neural Representations. Neural implicit representations, such as ONet [26], DeepSDF [30], and NeRF [27], parameterize a continuous occupancy/SDF/radiance fields with a network (e.g. a multilayer perceptrons, MLP). The property (e.g. occupancy, color/density) of any location in the scene could be queried by feeding its coordinate (optionally along with specified viewing direction) to the MLP, which is also referred as the coordinate network. Rapid progresses have been achieved on shape modeling [3, 39, 24, 36] and appearance modeling [33, 10].
@@
Later, these neural implicit representations are extended to handle dynamic scenes (e.g. a talking head) by introducing some deformation techniques. For example, NerFace [6] conditions the NeRF on 3DMM coefficients, making a small step to animatable avatar creation. DFA-NeRF [42] further conditions NeRF with disentangled face attribute features and enable more detailed control over the talking head.
@@
Methods using neural implicit representations assume no fixed topology and have infinite resolution (in theory), achieving impressive novel view synthesis results for dynamic scenes. However, they usually suffer from various artifacts (e.g. blur, distortion) and inaccurate control when rendering novel expressions/poses, which is critical for creating high-quality animatable avatars. In contrast, our approach employs an explicit point-based representation driven by a coarse surface (represented as a UV position map) of the novel expression, achieving better generalization on novel expression and precise expression control.
@@
Discrete Neural Representations. Common discrete representations in CV/CG include mesh, regular grid/voxel, and irregular point cloud. Recently, these representations have been extended with neural features to not only reconstruct the shape, but also create photorealistic image renderings thanks to the rapid progress in differentiable neural rendering [31, 13, 32, 14].
@@
For example, neural volumes [20] learns a radiance field in the canonical space and introduce another warp field to handle dynamic scenes. However, regular grid/voxel-based representations usually suffer from the required cubic memory footprint and slow rendering speed due to the processing of empty voxels.
@@
Another line of works use mesh-based neural representations [21, 23]. For example, PiCA employs neural UV texture map to render the head avatar in different expressions from any given view point. Neural head avatar [7] introduces additional geometry and texture networks in complementary to a base FLAME [15] head model, resulting in better shape and texture modeling. However, mesh-based representations require accurate surface geometry and semantically consistent correspondence, which is usually hard to acquire. Alternatively, MVP combines mesh and grid-based representations, and uses the volume rendering technique to rendering images. But the color and density of MVP’s primitives are decoded from 2D CNNs and further aggregated to obtain the radiance of query points, which inevitably causes blurry renderings. As a result, we propose to explore the more flexible neural point-based representation. Concurrent to our work, PointAvatar [44] also uses a point-based representation. Different from ours, their points only store color information and use splatting for rendering. We utilize powerful volume rendering and thus can naturally render hairs in higher quality.
@@
Differentiable Neural Rendering. Rapid progress in differentiable neural rendering [9, 17, 18] makes analysis-by-synthesis pipeline more powerful than before and enables avatar creation from even a short monocular video [44, 7]. For example, deferred neural rendering [35] proposes a neural texture map, which is decoded by a neural renderer into high quality renderings from any viewpoint. Recently, neural volume rendering [27] achieves impressive high-quality renderings by introducing volume rendering to neural implicit fields. What’s more, volume rendering can naturally model translucent objects (e.g. hair, smoke) in very high fidelity, which is very suitable for head modeling. However, a major drawback of neural volume rendering is low rendering efficiency. Our approach employs volume rendering to ensure high-fidelity visual results, and propose a series of acceleration methods to achieve rendering speed matching neural texture rendering.
3 Methods
@@
In Sec. 3.1, we introduce our NPVA representation that leverages flexible point clouds for improved modeling of topological changes and thin structures, and utilizes volume rendering to produce high-fidelity images. In Sec. 3.2, we present our efficient rendering and training strategies, enabling NPVA to render photorealistic images comparable to NeRF while being 70$\sim$ 70$\times$ 70 × faster. Sec. 3.3 details our training losses. Sec. 3.4 describes our network architectures and implementation details.
3.1 Neural Point-based Volumetric Avatar
3.1.1 Neural Radiance Field
NeRF [27] effectively encodes a static scene using a Multi-Layer Perceptron network (MLP), achieving unprecedented high-quality novel view synthesis results. During rendering, the trained MLP takes the scene coordinates $\bm{x}\in\mathbb{R}^{3}$ and viewing direction $(\theta,\phi)$ as input and produces the corresponding density $\sigma_{x}$ and view-dependent color $\bm{c}_{\bm{x}}$. The final color of an image pixel is obtained via volume rendering (Eq. (1)), which integrates all shading points on the ray that passes through this image pixel. Mathematically, NeRF uses piece-wise constant density and color as an approximation:
$c=\sum^{N}_{i=1}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))c_{i},T_{i}=\exp(-\sum^{i-1}_{j=1}\sigma_{j}\delta_{j})$ (1)
@@ 3.1.1 Neural Radiance Field
3.1.2 Animatable Neural Points
@@
We employ explicit neural points in our NPVA to achieve more controllable deformation of the underlying implicit neural radiance field. Leveraging this geometry proxy can also significantly enhances the rendering efficiency (Sec. 3.2).
@@
Inspired by Point-NeRF [41], our representation consists of a set of neural points, denoted as $\mathcal{A}=\{(\bm{p}_{i},\bm{f}_{i})|i=1,\ldots,N\}$, where $\bm{p}_{i}\in\mathbb{R}^{3}$ denotes the location of point $i$, and $\bm{f}_{i}$ represents its associated feature. In our network, these features are learned in a variational auto-encoding style, similar to DAM [19] and PiCA [23], to create an animatable avatar. Specifically, the geometry sub-network in the decoder predicts a UV position map $\hat{\bm{G}}_{o}$ and a UV displacement map $\hat{\bm{G}}_{d}$. $\hat{\bm{G}}_{o}$ stores the vertex positions of a coarse head mesh. We apply additional intermediate supervision on the $256^{2}$ position map $\hat{\bm{G}}_{o}$ to obtain better expression control. To achieve better expressiveness of the point-based neural radiance field, we upsample the position map $\hat{\bm{G}}_{o}$ to $1024^{2}$ and incorporate the high-resolution displacement map $\hat{\bm{G}}_{d}$ to compensate for inaccuracy contained in the coarse geometry. By compositing the upsampled $\hat{\bm{G}}_{o}$ and $\hat{\bm{G}}_{d}$, we determine the positions of the neural points. The final neural points can adjust their positions adaptively around the surface, as we only apply a regularization term to penalize unreasonably large displacements.
@@
Figure 3: Lightweight Radiance Decoding. Given a query point, we find its $K$ nearest neighboring neural points. We weighted sum these points to obtain an “average” feature for the subsequent radiance decoding. Note that we remove the per-point processing MLP used in Point-NeRF. This lightweight decoding process runs faster and obtains better generalization on novel expressions for our dynamic modeling task.
3.1.3 Lightweight Radiance Decoding

### v3

@@ Abstract.
Rendering photorealistic and dynamically moving human heads is crucial for ensuring a pleasant and immersive experience in AR/VR and video conferencing applications. However, existing methods often struggle to model challenging facial regions (e.g., mouth interior, eyes, and beard), resulting in unrealistic and blurry results. In this paper, we propose Neural Point-based Volumetric Avatar (NPVA), a method that adopts the neural point representation as well as the neural volume rendering process and discards the predefined connectivity and hard correspondence imposed by mesh-based approaches. Specifically, the neural points are strategically constrained around the surface of the target expression via a high-resolution UV displacement map, achieving increased modeling capacity and more accurate control. We introduce three technical innovations to improve the rendering and training efficiency: a patch-wise depth-guided (shading point) sampling strategy, a lightweight radiance decoding process, and a Grid-Error-Patch (GEP) ray sampling strategy during training. By design, our NPVA is better equipped to handle topologically changing regions and thin structures while also ensuring accurate expression control when animating avatars. Experiments conducted on three subjects from the Multiface dataset demonstrate the effectiveness of our designs, outperforming previous state-of-the-art methods, especially in handling challenging facial regions.
Neural representation, volume rendering, high-fidelity head avatars
submissionid: 612journalyear: 2023copyright: rightsretainedconference: SIGGRAPH Asia 2023 Conference Papers; December 12–15, 2023; Sydney, NSW, Australiabooktitle: SIGGRAPH Asia 2023 Conference Papers (SA Conference Papers ’23), December 12–15, 2023, Sydney, NSW, Australiadoi: 10.1145/3610548.3618204isbn: 979-8-4007-0315-7/23/12ccs: Computing methodologies Rendering
1. Introduction
@@
Realizing photorealistic rendering of an animatable human head is a pivotal goal in computer graphics and vision, which has broad applications such as AR/VR communications (He et al., 2020; Orts-Escolano et al., 2016; Lombardi et al., 2018), gaming (Waggoner, 2009), and remote collaboration (Wang et al., 2021a). However, providing a satisfying and immersive experience in these applications remains immensely challenging due to our innate ability to express and perceive emotions through subtle facial cues (Ekman, 1980). Existing data-driven learning methods often generate noticeable artifacts in the mouth area and blurry beard textures (Zollhöfer et al., 2018; Egger et al., 2020; Khakhulin et al., 2022; Grassal et al., 2022; Lombardi et al., 2018; Ma et al., 2021). This limitation is primarily attributed to their underlying mesh-based representations, since the predefined mesh has a fixed topology and limited discretization resolution.
@@
To illustrate this limitation, consider the Deep Appearance Model (DAM) (Lombardi et al., 2018) that decodes a head mesh and the corresponding view-specific texture UV map. Despite achieving high-quality rendering results for the skin regions, DAM produces conspicuous artifacts in the mouth and hair regions due to inaccurate correspondences across frames (e.g., mouth interior) and the mesh’s inability to model thin structures (e.g., beard). To alleviate these issues, Pixel Codec Avatar (PiCA) (Ma et al., 2021) proposes the use of neural textures, allowing the subsequent neural renderer to address inaccurate shape estimation and topological inconsistencies (e.g., closed/open mouth), attaining moderately improved facial geometry and renditions for the mouth region. Similarly, Mixture of Volumetric Primitives (MVP) (Lombardi et al., 2021) attaches volumetric primitives (predicted by a CNN) to mesh vertices, replacing vertex colors with more flexible volumetric primitives. However, these enhanced “texture”-like representations remain embedded in a predefined topology and tend to produce blurry results if inaccurate correspondences occur.
@@
Therefore, we propose Neural Point-based Volumetric Avatar (NPVA) that leverages highly flexible neural points (Aliev et al., 2020) and versatile neural volume rendering, enabling sharper rendering for both topologically changing geometries (e.g., mouth interior) and translucent thin beard structures.
@@
Efficient rendering and training are also indispensable for practical applications. To this end, we propose three technical innovations. (1) We introduce a novel depth-guided (Lin et al., 2022) sampling method that incorporates local depth context information (i.e. a patch), achieving more realistic rendering while reducing the rendering time by $\sim 10\times$ compared to the vanilla NeRF. (2) We develop a lightweight radiance decoding process that eliminates unnecessary per-point processing used in (Xu et al., 2022; Bai et al., 2023), significantly improving rendering efficiency ($\sim 7\times$) and offering better generalization in our dynamic modeling task. (3) Lastly, to speed up training, we propose a novel Grid-Error-Patch (GEP) ray sampling strategy comprising three stages: a uniform grid-sampling stage for rapid initialization of a coarse result, an error-based importance sampling that delves into more challenging regions, and a patch-based stage to impose high-level perceptual image losses.
@@
  • We propose a novel volumetric representation based on neural points that are dynamically allocated around the surface (i.e., the target expression) for animatable head avatar creation. This representation is inherently capable of better handling thin geometry and topological changes.
@@
  • Experiments on the Multiface dataset show that our approach produces higher-quality images for novel expressions and novel views while being $\sim 70\times$ faster than NeRF.
@@
Figure 2. Overview of Neural Point-based Volumetric Avatar (NPVA). The core of our approach is a neural point-based volumetric representation (middle), with points distributed around the surface of the target expression. This surface is defined by the low-resolution position map $\hat{\bm{G}}_{o}$ with intermediate supervision. A high-resolution displacement map $\hat{\bm{G}}_{d}$ allows the points to adaptively move within a certain range, as needed to provide increased capacity in more challenging regions (e.g., mouth, hair/beard). The attached point features are obtained from the feature map $\hat{\bm{F}}$. $\hat{\bm{G}}_{o}$, $\hat{\bm{G}}_{d}$, and $\hat{\bm{F}}$ are decoded from the latent code $\bm{z}$ (left), which is trained in a variational auto-encoding style (encoder omitted). In addition, we propose three technical innovations with the aim of achieving rendering efficiency on par with mesh-based methods (right).
2. Related Works
2.1. Continuous Neural Representations.
@@ 2. Related Works
Neural implicit representations, such as ONet (Mescheder et al., 2019), DeepSDF (Park et al., 2019), and NeRF (Mildenhall et al., 2020), parameterize a continuous occupancy/SDF/radiance fields with a network (e.g., a multilayer perceptrons, MLP). The property (e.g., occupancy, color/density) of any location in the scene could be queried by feeding its coordinate (optionally along with specified viewing direction) to the MLP, which is also referred as the coordinate network.
@@
Rapid progresses have been achieved on shape modeling (Chen and Zhang, 2019; Wang et al., 2021b; Martel et al., 2021; Tretschk et al., 2020; Wu et al., 2023b) and appearance modeling (Sitzmann et al., 2019; Schwarz et al., 2020; Kellnhofer et al., 2021; Zheng et al., 2023b). For example, PVA (Raj et al., 2021) and KeypointNeRF (Mihajlovic et al., 2022) obtain facial neural radiance fields from multi-view images for high-fidelity static face reconstruction. Later, these neural implicit representations are further extended to handle dynamic scenes (e.g., a talking head) by introducing deformation techniques (Yu et al., 2023; Gafni et al., 2021; Yao et al., 2022; Tang et al., 2022; Wu et al., 2023a). For instance, NerFace (Gafni et al., 2021) conditions the NeRF on 3DMM coefficients, making a small step to animatable avatar creation. DFA-NeRF (Yao et al., 2022) further conditions NeRF with disentangled face attribute features and enables more detailed control over the talking head.
@@
Methods using neural implicit representations assume no fixed topology and have infinite resolution (in theory), achieving impressive novel view synthesis results for dynamic scenes. However, they usually suffer from various artifacts (e.g., blur, distortion) and inaccurate control when rendering novel expressions/poses, which is critical for creating high-quality animatable avatars. In contrast, our approach employs an explicit point-based representation driven by a coarse surface (represented as a UV position map) of the novel expression, achieving better generalization on novel expression and precise expression control.
2.2. Discrete Neural Representations.
@@
Common discrete representations in CV/CG include mesh, grid/voxel, and point cloud. Recently, these representations have been extended with neural features to not only reconstruct the shape, but also create photorealistic image renderings thanks to the rapid progress in differentiable neural rendering (Rakhimov et al., 2022; Laine et al., 2020; Rückert et al., 2022; Lassner and Zollhöfer, 2021; Cao et al., 2022; Ma et al., 2023; Wang et al., 2022, 2021c).
@@
For example, neural volumes (Lombardi et al., 2019) learns a radiance field in the canonical space and introduce another warp field to handle dynamic scenes. However, regular grid/voxel-based representations usually suffer from the required cubic memory footprint and slow rendering speed due to the processing of empty voxels.
@@
Another line of works uses mesh-based neural representations (Lombardi et al., 2021; Ma et al., 2021; Wang et al., 2025, 2023). For example, PiCA employs a neural UV texture map to render the head avatar in different expressions from any given viewpoint. Neural head avatar (Grassal et al., 2022) introduces additional geometry and texture networks in complementary to a base FLAME (Li et al., 2017) head model, resulting in a better shape and texture modeling. However, mesh-based representations require accurate surface geometry and semantically consistent correspondence, which is usually hard to acquire. Alternatively, MVP combines mesh and grid-based representations, and uses the volume rendering technique to render images. However, the color and density of MVP’s primitives are decoded from 2D CNNs and further aggregated to obtain the radiance of query points, which inevitably causes blurry renderings. As a result, we propose to explore a more flexible neural point-based representation. Concurrent with our work, PointAvatar (Zheng et al., 2023a) also uses a point-based representation. Different from ours, their points only store color information and use splatting for rendering. We utilize powerful volume rendering, which has the potential to render higher-quality hairs and beards.
2.3. Differentiable Neural Rendering.
@@
Rapid progress in differentiable neural rendering (Jiang et al., 2020; Liu et al., 2019, 2020; Müller et al., 2022; Wang et al., 2019; Sun et al., 2025) makes analysis-by-synthesis pipeline more powerful than before and enables avatar creation from even a short monocular video (Zheng et al., 2023a; Grassal et al., 2022). For example, deferred neural rendering (Thies et al., 2019) proposes a neural texture map, which is decoded by a neural renderer into high quality renderings from any viewpoint. Recently, neural volume rendering (Mildenhall et al., 2020) achieves impressive high-quality renderings by introducing volume rendering to neural implicit fields. What’s more, volume rendering can naturally model translucent objects (e.g., hair, smoke) in very high fidelity, which is very suitable for head modeling. However, a major drawback of neural volume rendering is low rendering efficiency. Our approach employs volume rendering to ensure high-fidelity visual results, and proposes a series of acceleration methods to achieve rendering speed matching neural texture rendering.
3. Methods
@@
In Sec. 3.1, we introduce our NPVA representation that leverages flexible point clouds for improved modeling of topological changes and thin structures, and utilizes volume rendering to produce high-fidelity images. In Sec. 3.2, we present our efficient rendering and training strategies, enabling NPVA to render photorealistic images comparable to NeRF while being $\sim 70\times$ faster. Sec. 3.3 details our training losses. Sec. 3.4 describes our implementation details.
3.1. Neural Point-based Volumetric Avatar
3.1.1. Neural Radiance Field
NeRF (Mildenhall et al., 2020) effectively encodes a static scene using a Multi-Layer Perceptron network (MLP), achieving unprecedented high-quality novel view synthesis results. During rendering, the trained MLP takes the scene coordinates $\bm{x}\in\mathbb{R}^{3}$ and viewing direction $(\theta,\phi)$ as input and produces the corresponding density $\sigma_{x}$ and view-dependent color $\bm{c}_{\bm{x}}$. The final color of an image pixel is obtained via volume rendering (Eq. (1)), which integrates all shading points on the ray that passes through this image pixel. Mathematically, NeRF uses piece-wise constant density and color as an approximation:
(1) $c=\sum^{N}_{i=1}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))c_{i},\;\;T_{i}=\exp(-\sum^{i-1}_{j=1}\sigma_{j}\delta_{j})$
@@ 3.1.1. Neural Radiance Field
3.1.2. Animatable Neural Points
@@
We employ explicit neural points in our NPVA to achieve more controllable deformation of the underlying implicit neural radiance field. Leveraging this geometry proxy can also significantly enhance the rendering efficiency (Sec. 3.2).
@@
Inspired by Point-NeRF (Xu et al., 2022), our representation consists of a set of neural points, denoted as $\mathcal{A}=\{(\bm{p}_{i},\bm{f}_{i})|i=1,\ldots,N\}$, where $\bm{p}_{i}\in\mathbb{R}^{3}$ denotes the location of point $i$, and $\bm{f}_{i}$ represents its associated feature. In our network, these features are learned in a variational auto-encoding style, similar to DAM (Lombardi et al., 2018) and PiCA (Ma et al., 2021), to create an animatable avatar. Specifically, the geometry sub-network in the decoder predicts a UV position map $\hat{\bm{G}}_{o}$ and a UV displacement map $\hat{\bm{G}}_{d}$. $\hat{\bm{G}}_{o}$ stores the vertex positions of a coarse head mesh. We apply additional intermediate supervision on the $256^{2}$ position map $\hat{\bm{G}}_{o}$ to obtain better expression control. To achieve better expressiveness of the point-based neural radiance field, we upsample the position map $\hat{\bm{G}}_{o}$ to $1024^{2}$ and incorporate the high-resolution displacement map $\hat{\bm{G}}_{d}$ to compensate for inaccuracy contained in the coarse geometry. By compositing the upsampled $\hat{\bm{G}}_{o}$ and $\hat{\bm{G}}_{d}$, we determine the positions of the neural points. The final neural points can adjust their positions adaptively around the surface, as we only apply a regularization term to penalize unreasonably large displacements.
@@
Figure 3. Lightweight Radiance Decoding. Given a query point, we find its $K$ nearest neighboring neural points. We weighted sum these points to obtain an “average” feature for the subsequent radiance decoding. Note that we removed the per-point processing MLP used in Point-NeRF. This lightweight decoding process runs faster and obtains better generalization on novel expressions for our dynamic modeling task.
3.1.3. Lightweight Radiance Decoding